# Uncertainty quantification in fine-tuned LLMs using LoRA ensembles

### v1

@@ 8 Conclusions
We derived a principled, and practically efficient method for posterior approximation of fine-tuned LLMs using ensembles of low-rank adaptation models. By analysing the evolution of entropic uncertainty measures during fine-tuning, we identified signals of dataset complexity and signs of architecture limitations.
The methods presented are generally applicable to different fine-tuning scenarios, and provide a systematic treatment of prediction uncertainty for fine-tuned LLMs.
9 Acknowledgements
HL was supported by a grant from the KAW Foundation. The computations were enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS) and the Swedish National Infrastructure for Computing (SNIC) at C3SE partially funded by the Swedish Research Council through grant agreements no. 2022-06725 and no. 2018-05973. OB would like to thank Ihor Balabanov for valuable discussions.
@@ A.1 Size
For MMLU STEM and Social Studies datasets, we merge the development and validation sets to create the training set and use the original test split as the validation set, as in (Wang et al., 2023). For CQA dataset we use the standard training and validation splits. The topics in the MMLU dataset are categorized into STEM, Social Studies, Humanities, and Other, following the classification defined in (Hendrycks et al., 2021).
@@ A.2 Examples
Q: A revolving door is convenient for two direction travel, but it also serves as a security measure at a what? Answer Choices: (a) bank (b) library (c) department store (d) mall (e) new york A: (a). CommonsenseQA
Q: Which one of the following is the most appropriate definition of a 99% confidence interval? Answer Choices: (a) 99% of the time in repeated samples, the interval would contain the true value of the parameter (b) 99% of the time in repeated samples, the interval would contain the estimated value of the parameter (c) 99% of the time in repeated samples, the null hypothesis will be rejected (d) 99% of the time in repeated samples, the null hypothesis will not be rejected when it was false A: (a). MMLU SS
Q: Find all c in Z_3 such that Z_3[x]/(x^2 + c) is a field. Answer Choices: (a) 0 (b) 1 (c) 2 (d) 3 A: (b). MMLU STEM

### v2

@@ 8 Conclusions
We derived a principled, and practically efficient method for posterior approximation of fine-tuned LLMs using ensembles of low-rank adaptation models. Using the posterior approximation, we derived uncertainty quantification using entropic metrics, and followed their evolution during fine-tuning. This enabled us to draw a number of conclusions regarding the balance of retained knowledge and domain specific adaptation. First, the low-rank ensemble shows retained accuracy while also decreasing both predictive uncertainty and mutual information in the overfitting regime. At the same time, incorrect predictions are increasingly correlated with higher mutual information, opening up for inference time filtering and active learning. Second, during fine-tuning towards CQA, there is a distinct difference between the change in uncertainty for MMLU STEM and MMLU social science. For the STEM dataset, there is a large fraction of samples with low mutual information and high predictive entropy, hinting at in-domain but uncertain predictions. This can be connected to the lower prior accuracy of MMLU STEM compared to MMLU SS and CQA, indicating low representative power of the model in this domain. These STEM samples exhibit increased robustness, through broad yet consistent predictive distributions across ensemble members, to a large extent surviving overfitting.
As fine-tuning continues to be an important tool in shaping model output Zhang et al. (2023b), it is also increasingly important to understand how the knowledge of the pre-trained model changes as the model is adapted to a target domain. The methods presented here are generally applicable to different fine-tuning scenarios, and provide a systematic treatment of prediction uncertainty for fine-tuned LLMs.
9 Acknowledgments
The computations were enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS) and the Swedish National Infrastructure for Computing (SNIC) at C3SE partially funded by the Swedish Research Council through grant agreements 2022-06725 and 2018-05973.
10 Impact Statement
This paper advances the field of Machine Learning by enhancing the reliability and understanding of large language models (LLMs) through uncertainty quantification using LoRA ensembles. Our work contributes to the understanding and trustworthiness of LLM predictions, a step forward in AI reliability and applicability.
@@ A.1 Size
For MMLU STEM and Social Studies datasets, we use the original test split as the validation set, as in Wang et al. (2023). For CQA dataset we use the standard training and validation splits. The topics in the MMLU dataset are categorized into STEM, Social Studies, Humanities, and Other, following the classification defined in Hendrycks et al. (2021).
@@ A.2 Examples
{mdframed}
[backgroundcolor=gray!20, linecolor=black, linewidth=1pt, roundcorner=10pt]
Q: A revolving door is convenient for two direction travel, but it also serves as a security measure at a what?
Answer Choices:
(a) bank
(b) library
(c) department store
(d) mall
(e) new york
A: (a).
CommonsenseQA
{mdframed}
[backgroundcolor=gray!20, linecolor=black, linewidth=1pt, roundcorner=10pt]
Q: Which one of the following is the most appropriate definition of a 99% confidence interval?
Answer Choices:
(a) 99% of the time in repeated samples, the interval would contain the true value of the parameter
(b) 99% of the time in repeated samples, the interval would contain the estimated value of the parameter
(c) 99% of the time in repeated samples, the null hypothesis will be rejected
(d) 99% of the time in repeated samples, the null hypothesis will not be rejected when it was false
A: (a).
MMLU SS
{mdframed}
[backgroundcolor=gray!20, linecolor=black, linewidth=1pt, roundcorner=10pt]
Q: Find all c in Z_3 such that Z_3[x]/(x^2 + c) is a field.
Answer Choices:
(a) 0
(b) 1
(c) 2
(d) 3
A: (b).
MMLU STEM